OpenStudy (kpop4life123):

FAST IM RUNNING OUT OF TIME !!!!!!! HELP ME!!!!!!!!!! I WILL FAN AND MEDAL !!!!!!! The sets of numbers 7, 24, 25 and 9, 40, 41 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples!

OpenStudy (astrophysics):

\[a^2+b^2=c^2\] and right triangles, that should suffice.

OpenStudy (anonymous):

Pythagorean triples is when all all the numbers of the pythagorean theorem are integers (whole numbers)

OpenStudy (unicornpoopcookies):

Well, one of the triples is 7, 24, 25. You need to show that 72+242=252

OpenStudy (kpop4life123):

So like is this like the answers or just helping ?

OpenStudy (anonymous):

if that's who I think it is, nice profile pic btw

OpenStudy (anonymous):

these are all hints, not answers

OpenStudy (kpop4life123):

gdragon?!

OpenStudy (kpop4life123):

lol

OpenStudy (anonymous):

lol that man is life...no exit

OpenStudy (anonymous):

But back to your question. I'll walk you through why set (7, 24, 25) is a pythagorean triplet

OpenStudy (unicornpoopcookies):

I'd give the answer if I knew it thats all I know :X

OpenStudy (unicornpoopcookies):

just give him or her the answer for Gods sake lol

OpenStudy (anonymous):

So we know that pythagorean triplets are when a, b and c are integers, meaning whole numbers.

OpenStudy (unicornpoopcookies):

RUNNING OUT OF TIME HERE

OpenStudy (kpop4life123):

i really dont know how to do this cuz i messed so much days of school cuz i was sick and im trying to hurry up cuz this is missed assignments and tomorrow is the last day

OpenStudy (unicornpoopcookies):

Its alright kpop,no ones gonna ask you about this theorum thing in real life.

OpenStudy (anonymous):

well just use the pythagorean theorem and prove that 7^2 + 24^2 = 25^ (as per the pythagorean theorem which is a^2 +b^2 = c^2)

OpenStudy (anonymous):

*25^2

OpenStudy (anonymous):

I keep repeating myself, but the difference between pythagorean triplets and a regular pythagorean theorem is that  for pythagorean triplets, a & b & c all have to be integers (meaning they must all be whole numbers)